### v1

# A Two-Stage Stochastic Model for Road-Rail Intermodal Freight Transportation Under Demand and Capacity Uncertainty

Abstract
With the steady increase in global logistics and freight transport demand, the need for efficient and sustainable intermodal transport systems becomes increasingly important. This study addresses the optimization of container movement by intermodal transport with fixed train schedules. We emphasize the integration of road-rail intermodal transport amid uncertain demand and train (spot) capacities. A two-stage stochastic optimization model is developed to strategically manage the transportation of containers from multiple origins to designated intermodal hubs. By leveraging spot capacities at train stations and addressing uncertainties in demand and train capacity, the model integrates Conditional Value-at-Risk (CVaR) to balance cost efficiency and risk management, enabling robust decision-making under uncertainty. The model’s objectives encompass minimizing transportation costs, mitigating carbon emissions, and enhancing the reliability of containerized freight movement across the network. A comprehensive case study using real-world data demonstrates the practical applicability of the model, highlighting its effectiveness in reducing operational costs, minimizing environmental impacts, and providing actionable insights for stakeholders to navigate the trade-offs between expected costs and risk management in dynamic intermodal transport settings.
Keywords Intermodal Freight transportation  $\cdot$ Stochastic Optimization  $\cdot$ Decarbonization  $\cdot$ Road-Rail Transport
1 Introduction
Intermodal freight transportation involves the movement of freight using at least two transportation modes without changing the load unit throughout the process. This approach leverages the cost-effectiveness and high capacity of modes like rail or sea for long-haul segments while relying on trucks for shorter distances. It typically requires consolidating multiple shipments into containers or single units at facilities such as warehouses or manufacturing points. These consolidated units are then transported to intermodal hubs for shipping and eventually transferred to final intermodal nodes for deconsolidation and final delivery. The scale and importance of intermodal freight are evident in the fact that, in 2021, the United States (US) alone transported an average of 53.6 million tons of freight daily, valued at over \$54 billion (Bureau of Transportation Statistics,, 2021), emphasizing it’s significant role in the national economy.
Despite its advantages, intermodal freight transport faces several challenges due to the involvement of multiple stakeholders and decision-makers, including shippers, carriers, and policymakers, each with their own priorities and objectives. In the context of our work, a significant challenge is the modal shift from road to rail transport (El Yaagoubi et al.,, 2022). Integrating these modes efficiently requires overcoming additional handling and time requirements compared to direct road transport (Görçün et al.,, 2024). For longer distances, rail’s high fixed costs are offset by increased efficiency, while road transport becomes less cost-effective as distance increases (Zgonc et al.,, 2019).
Synchronized operations within intermodal networks are very crucial, especially when integrating modes with fixed schedules, such as rail transport with more flexible modes like road transport. For instance, trucks must arrive at rail terminals in time to accommodate service time for modal shift and align with train departure schedules. Achieving this requires precise planning and real-time adjustments to account for potential disruptions such as traffic congestion, accidents, or weather-related delays. Effective coordination minimizes delays and ensures the overall efficiency of the intermodal transport system.
Another critical challenge is managing the uncertainty in market demand and spot capacities at intermodal hubs. Demand fluctuations can arise from various factors, including booking changes, cancellations, unforeseen events, and seasonal variations (Wang and Meng,, 2021). Compounding this uncertainty is the stochastic nature of spot capacities, which depend not only on competing cargo from other companies but also on the variability of loading and unloading activities at each station (Sun and Schonfeld,, 2016). These dynamics create a complex environment where maintaining operational efficiency requires careful planning and adaptability.
A key aspect influencing spot capacity allocation is the role of long-term contracts. These agreements are typically made between transport providers and large companies or freight forwarders, ensuring a guaranteed capacity over a specified period. Such contracts offer stable and predictable volumes for the transport providers while securing regular service for the customers. However, any unused capacity from these agreements is often redirected to the spot market. This cascading allocation process introduces additional uncertainty for spot customers, as the availability of capacity depends on how much of the allocated slots long-term contract holders utilize (Wang and Meng,, 2021). For instance, if contract holders underutilize their guaranteed slots, the surplus is released for spot demand, but this availability may still be constrained by the need to accommodate delayed or overbooked cargo from other services or other long-term contractors. This interplay between long-term contracts and spot demand underscores the need for flexible and adaptive strategies to ensure optimal utilization of available capacity while addressing the inherent uncertainty in both supply and demand (Lu et al.,, 2020; Delbart et al.,, 2021).
In this study, we address the problem faced by manufacturing and logistics companies in managing intermodal transport operations with predefined and fixed train schedules, specifically focusing on container transport. Our primary emphasis is on the critical modal shift from road to rail, highlighting the need for precise timing and coordination to meet fixed train schedules efficiently and reliably. Containers are transported from multiple warehouses or manufacturing plants to intermodal hubs or train stations using trucks. However, we do not consider movements beyond the final train stations, as these often serve as centralized distribution hubs with independent local logistics networks (Rodrigue,, 2020). By narrowing the scope to the intermodal portion of the supply chain, we aim to optimize this resource-intensive segment, which offers significant opportunities for efficiency and environmental benefits.
This problem is further complicated by the stochastic nature of demand and uncertainties in spot capacities at each station, as highlighted earlier. Companies must determine the optimal number of containers to prepare at supply nodes, select facilities for container transport, and identify suitable intermodal hubs or train stations for delivery, all while accounting for variability in spot capacity availability (see Fig. 1). The complexity is heightened by multiple trains with fixed schedules and the cascading effects of unused long-term contract capacities on spot market availability. Our approach incorporates sustainability objectives by minimizing carbon emissions during transportation, enhancing both operational efficiency and environmental performance.
Figure 1: Overview of the intermodal transport system: Demonstrating the strategic integration and modal shift between road and rail networks for container transport
We propose a two-stage stochastic optimization model that integrates carbon reduction strategies. The objective of the model is to minimize a mean-risk function that accounts for various components of the total random cost. These include the operational cost of preparing containers at the origins, transportation costs from origins to intermodal hubs or train stations, penalties for unmet demand, and penalties for exceeding carbon emissions limits. To address the uncertainty inherent in demand and available capacity at each station, our model incorporates Conditional Value-at-Risk (CVaR) as part of the objective function. CVaR is a widely-used risk measure that focuses on minimizing the potential for extreme losses. It accounts for the “tail risk" by capturing the expected losses exceeding a specified quantile in the upper tail of the loss distribution. In our case, CVaR helps balance cost efficiency with risk management by emphasizing the potential downside risks associated with demand fluctuations and capacity uncertainties.
In the first stage, the model determines the optimal supply levels at each origin and the initial transportation of containers to train stations before the uncertainty is realized. As for the second stage, after the uncertainty is realized, the model re-optimizes the allocation of containers to meet the demand as closely as possible utilizing the available capacity while minimizing penalties for unmet demand and excess emissions. The model ensures the transportation network adapts to actual conditions, balancing the trade-offs between demand and capacity requirements, costs, and operational efficiency under stochasticity.
To showcase the practical application of our model, we incorporate a case study derived from a segment of the extensive rail network in the US. This case study examines the integration of road and rail transportation, focusing on five critical freight train routes that link a total of 26 intermodal hubs with 68 FedEx warehouses distributed across multiple states, ensuring a minimum of two warehouses per state. Utilizing real-world data on train schedules, distances, and facility locations from the Freight Analysis Framework (FAF) and the National Transportation Atlas Database (NTAD), the study offers a comprehensive view of the operational dynamics within the network.
The remaining sections of this paper are organized as follows. Section 2 presents a literature review on intermodal transport resilience. Section 3 provides a detailed problem description. Section 4 discusses our proposed optimization model. Section 5 presents our computational experiments and insights from sensitivity analysis. Section 6 concludes the paper and suggests directions for future research.
2 Literature review
Intermodal freight transport has been widely studied for its potential to improve logistics efficiency, sustainability, and resilience. Recent advancements in optimization techniques, data-driven decision-making, and sustainable transport policies have addressed key challenges such as managing uncertainties, coordinating operations across modes, and minimizing emissions. This section highlights significant recent contributions in these areas, examines emerging approaches, and identifies research gaps, providing context for the proposed model’s contributions.
2.1 Managing Uncertainties in Intermodal Transport
Managing uncertainties in intermodal transport has been a focal point for researchers aiming to enhance the resilience and efficiency of logistics systems. Guo et al., (2020) tackled dynamic and stochastic shipment allocation challenges by proposing a multistage stochastic programming model. Their anticipatory optimization approach effectively integrated rolling horizon techniques to manage future demand fluctuations, outperforming traditional methods in reducing costs and improving allocation efficiency. Similarly, Sun, (2020) employed a fuzzy multi-objective routing model to address demand uncertainties in hazardous materials transportation. By incorporating triangular fuzzy numbers and soft time windows, their model balanced risk, economic considerations, and service reliability, providing a comprehensive solution for uncertainty in freight transport.
Facility location optimization and disruption management have also been critical to addressing transport uncertainties. Badyal et al., (2023) introduced a two-stage stochastic model to optimize terminal locations and freight distribution in networks prone to disruptions. Their findings emphasized the need to integrate robust planning to enhance cost and resilience. Extending this focus, Bagheri et al., (2024) compared deterministic, stochastic, and robust approaches for optimizing dry port locations. Their study highlighted the superior adaptability of robust models to dynamic conditions, showcasing their potential to reduce environmental impacts while maintaining operational efficiency.
Routing and network optimization remain pivotal in mitigating operational uncertainties. Lu et al., (2020) introduced a fuzzy programming model to manage intercontinental road-rail transport uncertainties, achieving improved efficiency through sensitivity analyses and defuzzification techniques. Yang et al., (2024) expanded on routing challenges by developing a bi-level genetic algorithm to optimize multimodal container transport paths, addressing time-sensitive and general cargo delivery constraints. This method proved effective in minimizing costs and ensuring timely deliveries, offering a scalable solution for complex logistics systems. Wang et al., (2024) complemented these efforts with a constrained optimization model designed to address marine transport delays, employing adaptive differential evolution to significantly reduce costs and delays in multimodal networks.
Integrating demand and capacity uncertainties within broader network designs has yielded impactful results. Li et al., (2023) proposed a multi-commodity flow model using Benders decomposition to optimize shared freight and passenger railway capacity, achieving a balance between operational costs and user satisfaction. Wang et al., (2022) introduced a tri-level programming model with a Kriging-surrogate-based algorithm to optimize large-scale multimodal networks. Their innovative approach demonstrated significant computational efficiency and scalability, solving real-world networks with over 120,000 origin-destination pairs.
Lastly, decision-making frameworks have played a key role in addressing systemic uncertainties and stakeholder priorities. Gandhi et al., (2024) employed a hybrid multi-criteria decision-making framework to evaluate modal shift risks and enablers, emphasizing the need for long-term partnerships and policy interventions. Raad and Rajendran, (2024) further contributed by developing a hybrid fuzzy stochastic model to design closed-loop dry port networks, enhancing resilience through scenario-based robust optimization. These frameworks underline the importance of integrating strategic decision-making tools to build sustainable and adaptive intermodal logistics networks.
2.2 Operational Synchronization in Road-Rail Intermodal Transport
Operational synchronization between road and rail is crucial for reducing delays and improving intermodal efficiency. Newman and Yano, (2000) pioneered train schedule optimization using decomposition techniques to streamline scheduling and meet delivery requirements. Kravets et al., (2021) introduced a multicriteria model to minimize transit times, demonstrating efficiency gains by consolidating cargo at rail stations. Liu et al., (2024) addressed intermodal port challenges with a flexible yard crane scheduling model, optimizing task sequencing and reducing truck-induced delays, particularly in high railway container ratio environments.
Scheduling and routing optimization have further contributed to improving the performance of intermodal networks. Chen et al., (2016) examined uncertainties in vehicle scheduling and dispatching at intermodal terminals, demonstrating significant improvements in logistics coordination and reduction in delay. Meanwhile, Chen et al., (2022) developed a bi-level optimization model integrating route planning and timetable adjustments, achieving high capacity utilization and reduced costs in China-Europe rail networks. Additionally, Tong et al., (2024) focused on emergency response scenarios, presenting a hybrid flow shop scheduling model and a memetic algorithm to minimize congestion and transshipment delays in road-rail multimodal systems. These studies underscore the importance of tailored scheduling solutions to address specific operational challenges in intermodal logistics.
Strategic planning and infrastructure improvements are essential for synchronized intermodal operations. Delgado et al., (2021) developed an optimization framework for decentralized terminal management, reducing costs and improving terminal planning in Portugal’s rail-road network. Li et al., (2022) tackled bottlenecks in China-Europe freight transport with a network equilibrium model that identified critical transfer points and guided infrastructure expansion. In emerging economies, Kumar and Anbanandam, (2020) highlighted poor infrastructure and limited transshipment technologies as major barriers, recommending targeted investments and policy interventions to enhance intermodal efficiency.
The resilience of intermodal systems to disruptions is another critical aspect of synchronization. Pasquale et al., (2022) proposed a two-stage modeling approach to analyze the ripple effects of disruptions, such as railway arc failures, on passenger and freight flows. Their simulations revealed key dependencies within intermodal networks, informing resilience strategies. Ke and Verma, (2021) extended this focus by developing a framework to manage disruption risks in rail-truck intermodal networks, employing predictive analytics to identify critical terminals and optimize recovery strategies. Additionally, Guo et al., (2024) developed International Multimodal Transport Connectivity (IMTC) index, which assessed transport performance and highlighted the resilience of the China Railway Express (CR Express) during COVID-19 disruptions. These findings underscore the importance of resilient design in ensuring seamless intermodal operations.
Finally, addressing systemic barriers and operational inefficiencies is essential for improving synchronization. Ke, (2022) explored robust optimization strategies for hazardous material transportation, focusing on rerouting, yard repairs, and third-party carrier utilization under yard disruptions. By enhancing network reliability and mitigating disruptions, their work highlighted the value of resilience-focused strategies. Similarly, Ahmady and Eftekhari Yeghaneh, (2022) developed a multi-objective optimization model that balances transport costs and system efficiency while addressing the criticality of infrastructure components, contributing to the broader goal of synchronized operations.
2.3 Decarbonization in Intermodal Freight Transport
The pursuit of decarbonization in intermodal freight transport has spurred innovative strategies leveraging optimization models and policy interventions. Cheng and Tan, (2019) developed a multimodal path optimization model that minimized carbon emissions and costs by incorporating carbon taxation and trading within time-constrained freight systems. Their genetic algorithm-based approach demonstrated the feasibility of achieving low-carbon transport strategies while maintaining operational timeliness. Similarly, Sun et al., (2023) introduced a fuzzy nonlinear programming model to optimize carbon-efficient road-rail intermodal routing, balancing environmental and operational goals through time-dependent and uncertain transport conditions.
Technological advancements and fuel alternatives have emerged as critical drivers for reducing emissions in intermodal transport. Hernandez et al., (2024) presented the NUFRIEND framework to evaluate alternative fuel technologies for freight rail decarbonization, focusing on biodiesel, e-fuels, hydrogen, and battery-electric systems. Their life-cycle and techno-economic analyses highlighted battery-electric locomotives as a viable solution, achieving a 46% reduction in emissions under specific cost conditions. Matsuyama et al., (2024) echoed the importance of technological innovation in Japan’s hinterland transport system, demonstrating that advanced technologies could drive emission reductions of up to 87%, surpassing the impacts of modal shifts.
Policy-driven modal shifts also play a pivotal role in reducing the carbon footprint of intermodal systems. Halim, (2023) evaluated the impact of rail freight enhancement policies in Indonesia, revealing that combined fiscal measures and infrastructure investments could increase rail freight’s modal share to 20% by 2030, leading to a 21.6% reduction in CO2 emissions. Similarly, Nassar et al., (2023) used a System Dynamics model to analyze the effectiveness of modal shift policies in Brazil, emphasizing the need for infrastructure investments alongside fiscal and regulatory measures to accelerate decarbonization efforts. These studies underline the necessity of integrating multimodal strategies with policy frameworks to achieve sustainable freight transport.
Efforts to align intermodal strategies with emission reduction goals have also focused on collaborative and game-theoretic approaches. Zhang et al., (2022) proposed a collaborative planning framework that optimized capacity utilization and reduced emissions by incorporating eco-label preferences in intermodal freight transport. In parallel, Tamannaei et al., (2021) used a Stackelberg game model to evaluate government interventions in road-to-rail modal shifts, finding that well-designed fuel taxes and subsidies could enhance sustainability metrics without significant economic drawbacks. These findings emphasize the value of coordinated efforts between stakeholders and policymakers to foster sustainable intermodal operations.
Incorporating consolidation strategies into intermodal networks has further enhanced decarbonization efforts. Goodarzi et al., (2024) proposed a two-stage stochastic model integrating environmental and social metrics to address disruption scenarios effectively. Their results demonstrated substantial cost savings and resilience improvements, highlighting the potential of consolidation investments to yield significant environmental and economic benefits. Similarly, López-Acevedo et al., (2024) showcased the feasibility of short-distance intermodal solutions, where strategically located terminals optimized rail usage, reducing emissions and transport costs in the Madrid region.
Together, these studies illustrate the multifaceted approaches required to decarbonize intermodal freight systems, emphasizing the importance of integrating technological advancements, policy-driven modal shifts, and strategic network optimizations. The collective insights from these works provide a robust foundation for sustainable freight transport planning that balances environmental and operational goals.
2.4 Our Contribution
Building on the advances in the literature, our work addresses critical gaps in managing uncertainties, operational synchronization, and decarbonization within intermodal freight transport systems. We propose a two-stage stochastic optimization model tailored for road-rail intermodal freight transport that addresses critical gaps in literature. Unlike previous works that focus on isolated aspects such as shipment allocation under demand uncertainties (Guo et al.,, 2020) or carbon-efficient routing (Sun et al.,, 2023), our model integrates these components within a unified framework. By leveraging CVaR to account for extreme cost scenarios, our approach provides a robust mechanism to manage uncertainties while balancing cost efficiency and environmental sustainability.
A key aspect of our work is its incorporation of real-time adaptive responses to stochastic elements such as train capacities and demand fluctuations through the utilization of spot capacities at intermodal hubs. This dynamic approach ensures that decisions regarding container preparation and allocation aligns with operational constraints while optimizing resource utilization to meet demand efficiently. In contrast to static models Badyal et al., (2023), which focus primarily on long-term facility location planning, our model actively integrates recourse actions, enabling responsive adjustments to narrow time windows and fluctuating capacities at intermodal hubs.
Another distinctive feature of our model is its dual emphasis on minimizing costs and reducing environmental impact. While prior studies such as Hernandez et al., (2024) have evaluated the potential of alternative fuels, our model embeds emissions constraints directly into the optimization framework. By penalizing excess emissions and incorporating these penalties into the objective function, we ensure that carbon footprint reduction is a central consideration, particularly for applications involving time-sensitive or high-stakes freight such as hazardous materials. This approach aligns with federal and state regulatory frameworks, such as the Clean Air Act, which mandates compliance with stringent greenhouse gas emission standards at the national level (Agency,, 2024), and California’s state-specific policy aiming to reduce emissions by 85% below 1990 levels by 2045 (Institute,, 2024). These regulations underscore the critical need to integrate emissions constraints into optimization models to ensure legal compliance while advancing broader environmental objectives.
Finally, the model’s ability to balance risk aversion and operational efficiency sets it apart. By combining expected cost minimization with CVaR as a risk measure, our approach provides decision-makers with the flexibility to accommodate diverse risk preferences while ensuring robust performance under varying scenarios. This integration of risk and cost efficiency underscores the model’s practical relevance for designing sustainable and resilient intermodal transport systems.
3 Problem Description
We address the optimization challenge faced by transportation and manufacturing companies managing intermodal logistics involving predefined and fixed train schedules. These entities operate multiple warehouses and/or manufacturing plants, with the goal of transporting containers from these facilities to intermodal or train stations in time for train departures. The utilization of spot capacities available on trains at intermodal hubs is particularly crucial in such operations. These capacities are often subject to significant randomness due to fluctuating demand and competing priorities from other users, such as long-term contracts and operational constraints (Wang and Meng,, 2021). Addressing these challenges requires robust planning and optimization strategies tailored to intermodal logistics.
The primary focus of this study involves intermodal transportation using two modes: rail (trains) and road (trucks). Containers must be transported by road from various supply locations to specific intermodal hubs, ensuring they meet the fixed schedules of departing trains. Each train operates on a predetermined route and schedule, with the company utilizing the available capacity to transport containers to their respective destinations. The demand for container transport is defined for each train, aligning with the company’s logistics requirements. Given that only one train typically serves a particular destination within a specified period, the optimization process focuses on allocating resources efficiently to meet these constraints.
A distinguishing feature of this problem is the inherent stochasticity of both demand and train capacities. Demand variability reflects real-world logistics scenarios, where requirements at destination nodes are often uncertain (Demir et al.,, 2016). Additionally, the exact train capacities available at intermodal hubs are unpredictable (Lu et al.,, 2020), as they are influenced by shared usage among multiple users, including long-term contract commitments and operational constraints. Similar challenges are observed in container shipping, where fluctuating demand and capacity uncertainties significantly impact operational decisions (Wang and Meng,, 2021; Nurminarsih et al.,, 2024). Addressing these uncertainties necessitates adaptive planning approaches capable of responding dynamically to real-time variations, ensuring both logistics reliability and cost-effectiveness.
The problem requires companies to first determine the number of containers to prepare at supply locations, such as warehouses or manufacturing plants, with an associated preparation cost. Containers are considered ready for shipment as-is, regardless of whether they are fully or partially loaded, once deemed ready for order fulfillment. The next step involves determining the optimal transportation plan, including the selection of supply locations and intermodal hubs, to minimize costs and meet demand at destination nodes within the specified time windows.
To make this discussion more relatable, consider a practical example from our case study. A freight train route from Seattle, WA, to Chicago, IL, involves intermediate stops at Spokane, WA; Billings, MT; Fargo, ND; and Minneapolis, MN. The company must decide which containers from nearby facilities should be sent to each intermediate hub to align with the train’s fixed schedule. Given limited capacity at each hub due to competing cargo, this decision involves balancing demand at the destination with the availability of containers at supply locations, all while accounting for potential delays and environmental considerations.
Key constraints include the fixed train schedules, stochastic capacity availability at each station, and the specified delivery time windows for shipments. Containers must be transported via road to align with train schedules, optimizing the assignment of shipments to trains while minimizing penalties incurred from unmet demand. Environmental considerations are also integral to the problem, with the objective of reducing carbon emissions by selecting the closest supply locations for shipments, thereby minimizing overall travel distances.
Routing decisions from the supply/origin locations to the intermodal hubs are excluded from this problem. This exclusion assumes that these trips are carried out frequently using pre-determined, established routes. For companies managing intermodal logistics, such routes are often fixed due to operational consistency and infrastructure constraints. Treating these routes as fixed simplifies the problem while remaining realistic for practical applications. However, incorporating dynamic routing considerations could be a valuable extension in scenarios where flexibility is necessary, such as changing traffic conditions or infrastructure disruptions. We leave this as a direction for future research.
By integrating these considerations, the problem focuses on designing a transportation plan that minimizes costs, aligns with operational constraints, and reduces environmental impact. The utilization of spot capacities at intermodal hubs ensures efficient resource allocation despite the inherent uncertainties in train schedules and capacities. This adaptive approach addresses real-world logistical complexities, making it applicable to diverse operational settings. We consider the following assumptions in our model:
  • Containers are shipped as-is once deemed ready for order fulfillment, regardless of whether they are fully or partially loaded, ensuring operational simplicity.
  • Supply locations are pre-loaded with containers ready for transport, and delays or costs associated with procuring empty containers are not considered.
  • Train routes, schedules, and travel times are fixed and known during the planning horizon. Manufacturing and logistics companies have no control over these decisions, as they are determined by rail operators and infrastructure constraints.
  • Traffic conditions and potential delays are incorporated into travel time estimates to enhance the robustness of the transportation plan.
  • Containers are uniform in size and weight, simplifying loading and transportation logistics.
  • Once loaded onto a train, containers remain onboard until they reach the designated demand point.
  • Routing decisions from supply locations to intermodal hubs are excluded, as these routes are assumed to follow established patterns typical of many-to-many distribution networks in intermodal transport, where OD pairs often utilize consistent and direct transport paths (Basallo-Triana et al.,, 2021).
This problem description highlights the complex interplay between operational constraints, environmental considerations, and uncertainty in intermodal transportation. Our structured optimization framework ensures adaptability to real-world logistics challenges by integrating spot capacities and demand uncertainty. The next section introduces the mathematical foundation of the model.
4 Optimization Model
This section presents a two-stage stochastic optimization model designed to address intermodal transportation challenges, particularly uncertainties in demand and train capacities. The model integrates a Conditional Value-at-Risk (CVaR) framework to balance cost efficiency and risk management. Key model components, including sets, parameters, and variables, are outlined in Table 1. The formulation is explained alongside a discussion of constraints, and a deterministic equivalent of the stochastic model is provided.
Table 1: Sets, Parameters, and Variables Definitions
Set Definition
$I$ Origin locations (warehouses/manufacturing centers)
$J$ Train stations/intermodal hubs
$T$ Time periods
$\Omega$ Scenarios
$N$ Trains
Parameter Definition
$C_{ij}$ Base transportation cost per container from origin $i$ to station $j$
$\tau_{ij}$ Travel time from origin $i$ to station $j$
$\alpha$ Confidence level for CVaR
$\lambda$ Trade-off coefficient between cost and risk
$\theta$ Value-at-Risk (VaR) threshold for cost
$\rho$ Penalty for excess emissions
$\pi_{\omega}$ Unit penalty cost for missing the train schedule under scenario $\omega$
$\epsilon$ Total allowable carbon emission for planning horizon
$E_{t}$ Carbon emission per container per time period $t$
$K_{jn\omega}$ Random capacity of train $n$ at station $j$ in scenario $\omega$
$O_{i}$ Operational cost for preparing containers at origin $i$
$p_{\omega}$ Probability of scenario $\omega$
$D_{n\omega}$ Demand for train $n$ in scenario $\omega$
$\sigma_{nj}$ Predefined schedule, specifying the time at which train $n$ will be at station $j$
$\kappa_{i}$ Maximum number of containers that can be prepared at origin $i$
$M$ A large positive constant
Variable Definition
$x_{ijn\omega t}$ Number of containers shipped from origin $i$ to station $j$ for train $n$ at time $t$ in scenario $\omega$
$y_{i}$ Number of containers prepared at origin $i$
$U_{n\omega}$ Unmet demand for train $n$ in scenario $\omega=\max(0,D_{n\omega}-\mathcal{I}_{j_{f}n\omega})$
$\eta_{\omega}$ Excess emissions in scenario $\omega$
$\mathcal{I}_{jn\omega}$ Inventory of containers on train $n$ at station $j$ in scenario $\omega$
$z_{ijnst}$ Binary variable indicating if a container is shipped from origin $i$ to station $j$ for train $n$ at time $t$ in scenario $\omega$
The two-stage structure is well-suited for handling stochastic factors in intermodal transportation. It enables informed initial decisions while accounting for second-stage adjustments based on realized demand and capacity constraints. This approach optimizes container preparation at supply nodes while deferring detailed allocation and fulfillment decisions until more information becomes available, improving adaptability in uncertain conditions.
Stochastic optimization has been widely applied in intermodal logistics for managing demand and capacity uncertainty. Previous studies Tolooie et al., (2020); Huang, (2021) have demonstrated the effectiveness of two-stage models in minimizing costs while responding to real-time disruptions. By leveraging this framework, our model ensures cost-efficient and resilient decision-making in dynamic logistics environments.
In our proposed model, we introduce a two-stage stochastic programming framework that integrates risk aversion. In this framework, both the objective function and the constraints in the second stage incorporate stochastic parameters in line with acceptable practice. Specifically, we adopt a mean-risk objective function that incorporates CVaR with $\alpha\in(0,1)$ being the confidence level. This choice of risk measure allows us to model the risk preferences of decision makers concerning the uncertain total cost, offering flexibility to accommodate various risk thresholds from no risk ($\alpha=0$) to worst-case scenarios ($\alpha\rightarrow 1$).
To ensure the robustness of our approach, we control unfulfilled orders using demand constraints by penalizing these violations in the objective function. The first-stage problem in the model determines the operational cost incurred for preparing the containers and the number of containers shipped from each supply location based on the capacity of each facility. The problem in the second-stage focuses on meeting the demand and utilizing the available capacity on each train at each station. From this perspective, our model aims to design a reliable network for logistics operations by imposing constraints that ensure demand is fulfilled with a high specified probability.
Let the first-stage decision vector be represented by $(a,b)$, the stochastic data vector $\zeta(\omega)$, and the second-stage problem with state $\gamma(\omega)$. The optimal stochastic objective value for the second-stage be denoted as $S^{\prime\prime}(a,b,\zeta(\omega),\gamma(\omega))$.
Given the discussion above the objective of the second stage problem is stated as:
$S^{\prime\prime}(a,b,\zeta(\omega),\gamma(\omega))=\min\left(\sum_{i\in I}\sum_{j\in J}\sum_{n\in N}\sum_{t\in T}C_{ij}x_{ijn\omega t}+\pi_{\omega}\sum_{n\in N}U_{n\omega}\right)$ (1)
The objective function in (1) aims to minimize total transportation cost from origins to train stations while also penalizing unmet demand for each train in every scenario. This dual-purpose objective ensures efficient logistics operations by optimizing transportation routes and mitigating the impact of any shortfalls in meeting the demand for each train in every scenario.
Choices are typically made before uncertainties are realized, with recourse decisions deferred to the second stage that addresses the stochasticity. Once the uncertainty is known, detailed second-stage distribution decisions can be addressed. Thus, first-stage logistics planning concentrates on the number of containers prepared at each facility, while the second-stage problem manages demand and capacity uncertainties. We now present the formulation for our two-stage stochastic road-rail intermodal transportation model.
4.1 Conditional Value-at-Risk (CVaR)
Before delving into the full stochastic optimization model, we provide an overview of the CVaR risk measure, which plays a critical role in its formulation. Initially introduced by Rockafellar et al., (2000) for financial portfolio optimization, CVaR has gained widespread adoption across various fields due to its ability to quantify and mitigate risks associated with extreme outcomes. For example, Künzi-Bay and Mayer, (2006) reformulated CVaR minimization problems as two-stage recourse models in stochastic programming, while Fábián, (2008) developed polyhedral representations to efficiently handle CVaR objectives and constraints in large-scale problems. For further insights into these foundational studies, see Rockafellar et al., (2000); Künzi-Bay and Mayer, (2006); Fábián, (2008).
CVaR measures the expected value of losses exceeding a specified Value-at-Risk (VaR) threshold, making it particularly suitable for capturing "tail risk" in uncertain scenarios. Based on the work of Rockafellar et al., (2000), CVaR can be mathematically expressed as:
$\text{CVaR}_{\alpha}(R)=\min\left\{\text{VaR}+\frac{1}{1-\alpha}\mathbb{E}([R-\text{VaR}]_{+})\mid\text{VaR}\in\mathbb{R}\right\}$ (2)
This holds for $R$ been a random variable, $\alpha\in(0,1)$ denotes the confidence level of the CVaR associated with the expected value of $R$ given that it surpasses the value of $\alpha$-quantile which is the threshold value of (2). The VaR threshold represents the $\alpha$-quantile of $R$ such that the probability of the loss exceeding this value is $1-\alpha$. The notation $[m]_{+}$ = $m\in\mathbb{R}_{+}$ captures non-negative values. To enhance computational efficiency, (2) can be reformulated as a linear program based on the work of Rockafellar et al., (2000) and further explained by Elçi and Noyan, (2018). We adopt this equivalent linear program formulation which is:
$\text{CVaR}_{\alpha}(R)=\min\left\{\theta+\frac{1}{1-\alpha}\sum_{\omega\in\Omega}p_{\omega}\pi_{\omega}\mid\pi_{\omega}\geq y_{\omega}-\theta,\;\forall\,\omega\in\Omega,\;\pi\in\mathbb{R}^{n}_{+},\;\theta\in\mathbb{R}_{+}\right\}$ (3)
CVaR is recognized as a coherent risk measure, satisfying fundamental principles such as monotonicity, subadditivity, and translation equivariance Artzner et al., (1999). It is also classified as a spectral risk measure (Acerbi,, 2002), providing an interpretation as a weighted average of adverse outcomes. The dual representation of CVaR highlights this characteristic:
[…]
This dual form presented in (4) illustrates the ability of CVaR to focus on the most extreme outcomes within the distribution of $R$, offering a robust framework for balancing risk and operational efficiency. Its practical relevance in intermodal transportation lies in its capacity to mitigate disruptions caused by demand variability and capacity uncertainties, as explored in the subsequent sections.
4.2 The intermodal stochastic optimization model
To address the complexities of managing intermodal logistics under uncertainty, we present a proposed model that integrates cost efficiency, risk mitigation, and environmental sustainability. This formulation emphasizes a dual-stage decision-making process, where initial resource allocation is followed by adaptive adjustments based on real-time conditions. By structuring the model in this way, it captures the dynamic interplay between strategic planning and operational flexibility, setting the foundation for the detailed objectives and constraints discussed below.
4.2.1 Objective
Min $\left(\sum_{i\in I}O_{i}y_{i}+(1-\lambda)\mathbb{E}\left(S^{\prime\prime}(a,b,\zeta(\omega),\gamma(\omega)))\right)+\lambda\text{CVaR}_{\alpha}\left(S^{\prime\prime}(a,b,\zeta(\omega),\gamma(\omega))\right)+\sum_{t\in T}\rho\eta_{\omega}\right)$ (5)
The objective function (5) integrates both the preparation (first-stage) and operational costs (second-stage) for the intermodal stochastic optimization model. This objective minimizes the mean-risk function related to the total stochastic cost, which includes:
  • The operational cost of preparing containers at origins, represented by ($\sum_{i\in I}O_{i}y_{i}$)
  • The expected transportation cost from origins to train stations, and penalties for unmet demand across all scenarios, $(1-\lambda)\mathbb{E}\left(S^{\prime\prime}(a,b,\zeta(\omega),\gamma(\omega))\right)$,
  • The CVaR at confidence level $\alpha$ of the random total cost, $\lambda\text{CVaR}_{\alpha}\left(S^{\prime\prime}(a,b,\zeta(\omega),\gamma(\omega)))\right)$ which accounts for extreme cost realizations,
  • Penalties for exceeding emissions limits, captured by $\sum_{t\in T}\rho\eta_{\omega}$.
This comprehensive objective ensures efficient logistics operations by balancing cost minimization, risk aversion, and sustainability goals, enabling adaptive decision-making under uncertain demand and train capacity.
4.2.2 Constraints
$\sum_{j\in J}\sum_{n\in N}\sum_{t\in T}x_{ijn\omega t}\leq y_{i},$ $\forall i\in I,\,\forall\omega\in\Omega$ (6)
$\sum_{i\in I}\sum_{t\in T}x_{ijn\omega t}\leq K_{jn\omega},$ $\forall j\in J,\,n\in N,\,\omega\in\Omega$ (7)
$x_{ijn\omega t}\leq Mz_{ijn\omega t},$ $\forall i\in I,\,j\in J,\,n\in N,\,t\in T,\,\omega\in\Omega$ (8)
$\sum_{i\in I}\sum_{j\in J}\sum_{n\in N}\sum_{t\in T}x_{ijn\omega t}E_{t}\tau_{ij}\leq\epsilon+\eta_{\omega},$ $\forall\omega\in\Omega$ (9)
$\mathcal{I}_{j_{1}n\omega}=\sum_{i\in I}\sum_{t\in T}x_{ij_{1}n\omega t},$ $\forall n\in N,\,\omega\in\Omega$ (10)
$\mathcal{I}_{jn\omega}=\mathcal{I}_{(j-1)n\omega}+\sum_{i\in I}\sum_{t\in T}x_{ijn\omega t},$ $\forall j\in J\setminus\{j_{1}\},\,n\in N,\,\omega\in\Omega$ (11)
$\sum_{i\in I}\sum_{j\in J}\sum_{t\in T}x_{ijn\omega t}+U_{n\omega}=D_{n\omega},$ $\forall n\in N,\omega\in\Omega$ (12)
$\mathcal{I}_{j_{f}n\omega}\leq D_{n\omega},$ $\forall n\in N,\,\forall\omega\in\Omega$ (13)
$\sum_{i\in I}x_{ijn\omega t}\leq y_{i},$ $\text{if }t+\tau_{ij}\leq\sigma_{nj}\quad\forall j\in J,n\in N,\omega\in\Omega,t\in T$ (14)
$\sum_{i\in I}x_{ijn\omega t}=0$ $\text{if }t+\tau_{ij}>\sigma_{nj}\quad\forall j\in J,n\in N,\omega\in\Omega,t\in T$ (15)
$y_{i}\leq\kappa_{i},$ $\forall i\in I$ (16)
$z_{ijn\omega t}\in\{0,1\},$ $\forall i\in I,\,j\in J,\,n\in N,\,t\in T,\,\omega\in\omega$ (17)
$x_{ijn\omega t},y_{i},\mathcal{I}_{jn\omega},U_{n\omega}\in\mathbb{Z}_{\geq 0},\quad\sigma_{nj},\eta_{\omega}\geq 0,$ $\forall i\in I,\,j\in J,\,n\in N,\,t\in T,\,\omega\in\Omega$ (18)
The constraints applied to the model effectively determine demand assignments, the number of containers on each train per station, and limit the carbon emissions during the planning period. To start with, constraints (6) ensure that the total number of containers transported from each origin $i$ does not exceed the available supply at the origin. Constraints (7) ensure that the number of containers transported to each train station does not exceed the available capacity on each train at a particular station. Further, constraints (8) ensure that container flows $x_{ijn\omega t}$ are linked to binary decision variables $z_{ijn\omega t}$. These constraints enforce that containers can only be shipped if the corresponding binary variable is activated ($z_{ijn\omega t}=1$). The large constant $M$ acts as an upper bound, preventing any flow $x_{ijn\omega t}$ unless the route is selected.
As for constraints (9), they ensure that the total emissions do not exceed the specified emissions cap plus any allowable excess. Constraints (10) initialize the inventory at the first station for each train and scenario, meaning the number of containers on the train at the first station will be the number of containers that were transported to that station. While constraints (11) update the inventory at each subsequent station based on the previous station’s inventory and the incoming flow of containers, constraints (12) ensure that the total number of containers transported to each train station and the unmet demand sum up to the total demand for each train. Constraints (13) ensure that the inventory at the final station $j_{f}$ for each train across all scenarios does not exceed the demand for that train. Constraints (14) dictate that containers from origin are shipped to a train station at a certain time only if the shipping time plus the travel time does not exceed the train’s scheduled departure time.
We make sure that no containers are transported from any origin to a station at a particular time if the container cannot meetup with the train schedule (see Constraints (15)). Constraints (16) ensure that the available supply at each origin $i$, does not exceed the maximum allowable supply at each origin. Lastly, the binary penalty constraints (17) and the non-negativity constraints (18) ensure logical consistency and enforce the non-negativity of the decision variables.
4.3 Deterministic equivalent of stochastic optimization model
To simplify implementation and analysis, the stochastic model is reformulated into its deterministic equivalent. This reformulation explicitly separates the CVaR term, improving the interpretability of risk-related costs and facilitating the identification of their individual contributions to the overall objective. Relating our stochastic formulation to the knapsack-style approach outlined in (4), we define the random total cost $R$ as:
$R=\left(\sum_{i\in I}O_{i}y_{i}+S^{\prime\prime}(a,b,\zeta(\omega),\gamma(\omega))+\rho\eta_{\omega}\right)$ (19)
where $S^{\prime\prime}(a,b,\zeta(\omega),\gamma(\omega))$ represents the optimal value of the second-stage problem for scenario $\omega$, and $\rho\eta_{\omega}$ captures the penalties for exceeding emissions limits.
In this context, the Value-at-Risk at confidence level $\alpha$ denoted $\text{VaR}_{\alpha}(R)$, establishes an upper threshold on the total cost that is exceeded with a probability of at most $1-\alpha$. However, $\text{CVaR}_{\alpha}(R)$ provides deeper insight by quantifying the expected total cost in scenarios where this threshold is surpassed. This measure captures the severity of extreme cost events, emphasizing potential losses beyond the VaR threshold.
Building on this risk assessment framework, we now present the reformulated deterministic equivalent of our stochastic model. The objective function in (5) becomes:
$\min$ $\sum_{i\in I}O_{i}y_{i}+(1-\lambda)\sum_{\omega\in\Omega}p_{\omega}\left(\sum_{i\in I}\sum_{j\in J}\sum_{n\in N}\sum_{t\in T}C_{ij}x_{ijn\omega t}+\pi_{\omega}\sum_{n\in N}U_{n\omega}\right)$
$+\lambda\text{CVaR}_{\alpha}\left(\sum_{i\in I}\sum_{j\in J}\sum_{n\in N}\sum_{t\in T}C_{ij}x_{ijn\omega t}+\pi_{\omega}\sum_{n\in N}U_{n\omega}\right)+\sum_{\omega\in\omega}\rho\eta_{\omega}$ (20)
Subject to:
$\quad\quad\quad(6)-(18)$ (21)
In addition to the reformulation above with objective (20), which explicitly includes CVaR, the model can be equivalently reformulated for clarity and highlighting separation of terms. This alternative formulation separates the CVaR component from the main summation of costs, providing a more interpretable representation of the objective. By isolating the CVaR term, it enhances the understanding of its individual contribution to the total cost and its role in risk management. This alternative form is particularly useful for both theoretical analysis and practical implementation, as it allows for a clearer distinction between the expected costs and the risk-related costs. To facilitate this understanding, we present the alternative objective function as follows: Alternatively, (20) can be written as:
$\min$ $\sum_{i\in I}O_{i}y_{i}+(1-\lambda)\sum_{\omega\in\Omega}p_{\omega}\left(\sum_{i\in I}\sum_{j\in J}\sum_{n\in N}\sum_{t\in T}C_{ij}x_{ijn\omega t}+\pi_{\omega}\sum_{n\in N}U_{n\omega}\right)$
$+\lambda\text{CVaR}+\sum_{t\in T}\rho\eta_{\omega}$ (22)
Also subject to (6) - (18) and the constraints below:
$\xi_{\omega}\geq c_{e\omega}-\theta,\quad$ $\forall\omega\in\Omega$ (23)
$\xi_{\omega}\geq 0,\quad$ $\forall\omega\in\Omega$ (24)
$\text{CVaR}=\theta+\frac{1}{1-\alpha}\sum_{\omega\in\Omega}p_{\omega}\xi_{\omega}$ (25)
Where $\xi_{\omega}$ denote the excess cost by which the cost exceeds the $\theta$ in each scenario and $c_{e\omega}$ represent the total transportation and penalty costs for every scenario. (23) ensures that the excess cost is calculated correctly based on the difference between the total cost in each scenario and the VaR. Constraint (24) ensures that excess is non-negative. While (24) defines CVaR as the VaR ($\theta$) plus the expected excess cost over all scenarios, scaled by $\frac{1}{1-\alpha}$. CVaR represents the expected value of the excess cost beyond the $\theta$, providing a measure of the average loss in the worst $(1-\alpha)\%$ of cases.
It is important to note that lower values of risk measures are preferred, aligning with the objective of minimizing potential extreme costs. The parameter $\lambda\in\{0,1\}$ serves as a weighting factor that balances the trade-off between the expected cost and the risk associated with cost variability. A higher $\lambda$ places more emphasis on risk aversion, potentially at the expense of higher expected costs.
In summary, the deterministic equivalent formulation provides a practical and computationally efficient framework that captures both the expected operational costs and the risks associated with uncertainties in demand and capacity. By explicitly incorporating Conditional Value-at-Risk (CVaR) into the objective function and constraints, the model empowers decision-makers to balance cost efficiency with risk mitigation, enabling the development of robust and resilient intermodal transportation plans. The next section illustrates the model’s applicability using real-world data from U.S. freight networks, highlighting its ability to address operational complexities effectively and offering practical insights through a detailed case study and computational experiments.
5 Case Study and Computational Results
This section applies the proposed two-stage stochastic optimization model to an intermodal transportation problem in the U.S. The analysis utilizes publicly available datasets, including FAF and NTAD, which provide detailed information on rail connections, intermodal terminals, and road links to hubs from supply locations such as warehouses and manufacturing facilities. To address data gaps, missing supply-hub connections were supplemented by manually retrieving data from Google Maps. Figures 2(a) and 2(b) provide a visual overview of the spatial layout of the U.S. rail network, highlighting connectivity between supply points, intermodal hubs, and key corridors. This foundational data supports the development of a network model, detailed subsequently.
(a) Overview of the rail tracks connections in the U.S.
(b) Intermodal hubs for freight rail in the U.S.
Figure 2: Rail infrastructure and intermodal hubs in the United States
The computational experiments were executed on a machine equipped with a 12th Gen Intel(R) Core(TM) i7-1265U processor, 16 GB RAM, and a 64-bit operating system. Each instance was solved independently using the Gurobi Optimizer 11.0.0, ensuring efficient handling of the model’s complexity. The implementation was performed in Python 3.11.4. The results of these experiments, which offer critical insights into the model’s behavior under real-world constraints, are detailed in the subsequent subsections.
5.1 Policy Questions
To guide the analysis and validate the applicability of the proposed model, the case study focuses on addressing the following research questions. These questions are designed to evaluate the model’s robustness, scalability, and practicality in handling real-world intermodal transportation challenges:
  1. 1.
    Balancing Cost, Risk, and Decision-Making Under Uncertainty: How does the model navigate trade-offs between cost efficiency and risk mitigation, particularly under varying values of the risk-aversion parameter and confidence level? How do these risk-based decisions affect cost structures, system resilience, and operational priorities across different stakeholder preferences? Additionally, how does risk aversion influence the stability and predictability of cost outcomes across multiple uncertainty scenarios?
  2. 2.
    Trade-Offs Between Cost and Sustainability: What are the trade-offs between cost efficiency and sustainability goals, particularly under varying levels of emissions constraints?
  3. 3.
    Adaptability to Uncertainty: How robustly does the model adapt to uncertainty in demand, capacity, and operational disruptions such as delays or unforeseen changes?
  4. 4.
    Scalability Across Networks: Can the model efficiently accommodate larger intermodal networks, including scenarios with additional supply nodes, hubs, and complex routing requirements?
  5. 5.
    Impact of Intermodal Transfers: To what extent do intermodal transfer times and associated costs influence key optimization metrics such as total cost, scheduling efficiency, and risk measures?
  6. 6.
    Impact of Capacity Constraints on Cost and Demand Satisfaction: How do train capacity limitations affect total system costs and unmet demand, and at what point do additional capacity increases yield diminishing returns?
  7. 7.
    Impact of Risk Preferences on Cost Stability and Scenario Variability: How does increasing $\lambda$ affect the variability of cost outcomes across different demand scenarios? At what point do additional risk-mitigation efforts yield diminishing returns in stabilizing total costs?
By addressing these questions, the case study evaluates the model’s ability to optimize intermodal logistics under uncertainty while balancing cost minimization, risk resilience, and sustainability objectives. The subsequent sections provide detailed analyses and insights into these aspects, demonstrating the model’s practical applicability and contributions to both research and real-world industry practices.
5.2 Case Study Setup
In this section, we detail the data used for our case study, which focuses on major freight train routes that connect 26 intermodal hubs with 50 supply locations, primarily warehouses. These routes were carefully selected to reflect the complexity of intermodal transportation in the U.S., capturing variation in demand, train capacity, and schedules. The train routes and intermodal hubs analyzed are predominantly operated by the Norfolk Southern Corporation, a major freight railroad company in the U.S. However, five of the hubs are jointly used by Norfolk Southern and Union Pacific Railroad, reflecting real-world infrastructure sharing. These shared hubs introduce coordination challenges such as scheduling and capacity constraints, which are addressed in the optimization model to enhance its applicability.
To ensure uniformity across the rail network, FedEx warehouses were used as the supply locations. This choice was made to standardize the analysis and maintain consistency in evaluating logistics across different regions. FedEx warehouses, known for their strategic distribution capabilities, provide a practical and representative basis for modeling supply points in intermodal transportation. Furthermore, some of these warehouses serviced multiple intermodal hubs, reflecting the flexibility and extensive reach of distribution networks in real-world logistics.
To further enhance the accuracy and applicability of the model, some parameter values were adopted from peer-reviewed works. For example, the planning time horizon was similar to the 60 hours (or 2.5 days) used in Ke and Verma, (2021). Additionally, Assadipour et al., (2016) estimated that approximately one hour is required to load, unload, or transfer an intermodal container, a value incorporated into the model to reflect realistic handling times at intermodal hubs. These parameter choices ensure the model aligns with observed industry practices and remains grounded in empirical evidence, balancing cost efficiency and operational feasibility.
5.2.1 Network Dynamics
Building on the case study setup, this section details the operational framework of the intermodal transportation network, focusing on train and road transportation dynamics. These elements are critical for understanding how the proposed model optimizes logistics under real-world constraints.
  • Train Operations: The train operations modeled in this study are based on fixed schedules with defined origin-destination pairs and intermediate stops. Below, we describe the operations of one representative train (Train A), which connects major industrial hubs and serves as a vital corridor for intermodal freight movement:
    • Route Description: Train A begins its journey in Buffalo, NY, and terminates in Detroit, MI (Delray), traversing key economic regions. Intermediate stops include Greencastle, PA; Columbus, OH; Cincinnati, OH; St. Louis, MO; and Decatur, IL.
    • Operational Characteristics: The total route covers approximately 1,642 miles, with each intermediate stop allocated six hours of dwell time to facilitate loading, unloading, and container transfers. These dwell times are integral to aligning train schedules with supply and road network connections.
    • Additional Routes: Other intermodal routes in the network are detailed in the appendix (Tables A.1, A.2, and A.3), reflecting diverse O-D pairs and operational constraints across the modeled system. These routes contribute to evaluating the scalability of the model.
  • Road Transportation: The road network forms a critical connection to the intermodal system. Each supply location is linked to its nearest intermodal hub based on a fixed distance metric, ensuring efficiency in routing. While dynamic routing is not considered in this study, this assumption reflects industry practices where companies rely on pre-determined routes for recurring freight movements. This fixed routing approach facilitates accurate cost and time estimates while maintaining alignment with real-world logistical operations.
5.3 Computational Experiment
This section presents a comprehensive analysis of the proposed model’s performance under various configurations and parameter settings addressing the proposed research questions. Through a series of computational experiments, we evaluate the sensitivity of key model outputs, such as total cost and risk measures, to changes in $\lambda$ (risk aversion weight) and $\alpha$ (risk confidence level). The experiments also examine the model’s ability to balance cost minimization with risk mitigation and assess its adaptability to different logistical scenarios.
To ensure clarity, several abbreviations are used throughout this section. OBJ denotes the stochastic model’s objective function, incorporating both first-stage and second-stage costs. E(TC) represents the expected total cost across all scenarios, while ASC refers to the average stochastic cost, capturing the mean realized costs across experiments. CVaR (Conditional Value-at-Risk) quantifies the expected cost of extreme scenarios beyond a specified threshold, determined by $\alpha$.
The risk-aversion parameter $\lambda$ controls the trade-off between expected cost minimization and risk aversion. It is varied from 0 to 1 in increments of 0.1 (i.e., $\lambda=0,0.1,0.2,\ldots,1$), allowing us to explore the full spectrum of decision-making priorities, from purely cost-focused strategies ($\lambda=0$) to highly risk-averse approaches ($\lambda=1$). Similarly, the risk confidence level $\alpha$ is tested at values of 0.25, 0.5, 0.75, and 1. These values provide insights into the model’s response to varying levels of risk tolerance, from focusing on extreme losses in a quarter of scenarios ($\alpha=0.25$) to considering worst-case outcomes ($\alpha=1$).
The demand scenarios analyzed in this study include extreme scenarios, where demand is significantly higher or lower than other scenarios, representing fluctuations that arises due to seasonal shifts, economic disruptions, or supply chain imbalances. These variations enable an assessment of system performance under both exceptionally high and low freight volumes, capturing intermodal hubs and train stations with the highest stochasticity and examining their impact on capacity uncertainty at each station. By incorporating these extreme cases, the model captures a realistic spectrum of uncertainty, ensuring that optimization strategies remain effective across a wide range of operating conditions.
By analyzing these scenarios, we aim to demonstrate the robustness and practical applicability of the model in optimizing intermodal logistics under uncertainty. The subsequent subsections provide a detailed breakdown of the experiments, starting with an analysis of the impact of $\alpha$ and $\lambda$ on the expectation and CVaR of the total cost.
5.3.1 Impact of Risk-Aversion Parameter and Confidence Level on Cost and CVaR
The interplay between $\lambda$ and $\alpha$ provides valuable insights into how the model balances total cost minimization and risk mitigation. Understanding how these parameters influences cost and risk is crucial in decision-making under uncertainty. The parameter $\lambda$ dictates the weight placed on the CVaR, effectively controlling the trade-off between minimizing average costs and reducing exposure to extreme cost scenarios. Meanwhile, $\alpha$ determines the quantile of the cost distribution used for CVaR calculations, thereby capturing the confidence level for considering risk reflecting the level of conservatism in risk management.
The combined effect of these parameters is presented in Table 2, summarizing the relationship between $\lambda$, $\alpha$, and key cost metrics, and further illustrated in Figures 3(a), 3(b), and 4. This analysis directly addresses the first policy question on balancing cost and risk by demonstrating how changes in $\lambda$ and $\alpha$ influence total cost and extreme-event mitigation. Additionally, it provides insights into how risk-based decision-making affects cost structures and system resilience.
Table 2: Results for different $\alpha$ and $\lambda$ values (scaled to $10^{3}$).
$\alpha=0.25$ $\alpha=0.5$
Lambda OBJ ASC E(TC) VaR CVaR OBJ ASC E(TC) VaR CVaR
0.0 238.2 210.1 226.5 150.0 256.0 238.2 210.1 226.5 167.7 297.6
0.1 241.1 209.3 225.7 149.8 255.0 244.6 211.1 227.3 176.2 293.7
0.2 244.0 210.9 225.9 150.6 254.8 251.7 214.5 229.2 176.2 293.4
0.3 246.7 210.5 226.4 150.2 255.9 257.3 212.9 228.3 176.2 293.7
0.4 248.6 214.1 227.4 150.2 255.7 264.2 212.1 226.9 174.2 292.0
0.5 251.6 210.7 225.8 150.6 254.7 270.2 218.7 229.8 177.6 291.5
0.6 253.5 216.7 229.8 160.2 256.2 276.0 220.3 232.7 178.6 296.1
0.7 255.7 218.2 230.2 162.4 256.1 281.9 220.5 232.9 178.3 296.2
0.8 258.2 216.0 229.6 155.6 257.2 286.5 225.8 236.1 189.5 294.0
0.9 260.8 218.2 230.1 164.2 255.6 292.4 223.2 234.7 185.7 294.0
1.0 263.2 226.6 234.0 173.8 254.7 297.2 240.0 245.1 202.8 291.9
$\alpha=0.75$ $\alpha=0.9$
Lambda OBJ ASC E(TC) VaR CVaR OBJ ASC E(TC) VaR CVaR
0.0 238.2 210.1 226.5 200.7 385.7 238.2 210.1 226.5 200.7 509.0
0.1 246.8 212.7 228.3 200.1 385.2 248.2 211.1 227.1 194.4 519.6
0.2 255.5 216.1 229.6 200.9 384.7 257.0 215.7 229.6 200.3 508.9
0.3 264.6 211.6 228.3 215.1 371.4 266.3 211.0 227.7 201.4 500.9
0.4 272.7 221.7 232.3 216.1 371.6 275.8 216.4 230.9 216.1 472.4
0.5 280.8 219.4 231.8 216.1 370.4 284.5 218.2 231.1 216.1 472.4
0.6 288.3 219.7 233.2 207.1 376.1 292.4 220.1 234.0 216.3 472.2
0.7 295.7 221.6 234.5 214.8 371.3 301.1 221.2 234.6 215.1 474.5
0.8 302.6 236.7 246.6 257.4 342.3 308.7 232.0 243.3 257.1 398.1
0.9 308.8 239.7 252.7 257.1 341.6 315.3 240.1 252.8 257.1 398.3
1.0 315.1 282.3 285.9 298.2 314.1 321.4 279.9 284.2 295.2 330.3
Trade-Off Between Cost and Risk Mitigation
Table 2 and Figure 3(a) illustrate the impact of increasing $\lambda$ on the objective value. As expected, higher $\lambda$ values result in a steady increase in the objective function across all $\alpha$ levels. This outcome is intuitive since a greater emphasis on extreme scenarios leads to higher overall costs, as the model prioritizes robust solutions that minimize worst-case losses. However, the rate of increase in cost varies with $\alpha$, with higher confidence levels (e.g., $\alpha=0.90$) showing a steeper rise compared to lower ones (e.g., $\alpha=0.25$). This suggests that risk-averse settings amplify the effect of risk mitigation efforts, making them increasingly expensive as confidence levels increases.
Conversely, Table 2 and Figure 3(b) show the diminishing CVaR values as $\lambda$ increases. The reduction in CVaR demonstrates the model’s ability to shift resources toward solutions that mitigate extreme cost realizations. However, beyond a certain point ($\lambda\approx 0.5$), additional increases in $\lambda$ yield only marginal improvements in risk reduction, indicating diminishing returns in CVaR minimization. Notably, this effect is more pronounced for higher $\alpha$ values, reinforcing the idea that extreme risk management strategies become increasingly costly and less effective beyond a specific risk-aversion threshold.
Sensitivity of Objective Value to Confidence Level ($\alpha$)
Table 2 and Figure 4 examine how varying $\alpha$ affects the objective function for fixed $\lambda$ values. Across all $\lambda$ settings, higher $\alpha$ values lead to increased objective values, reflecting the model’s stronger emphasis on extreme costs as confidence levels rise. However, this effect is particularly sensitive to $\lambda$, with higher values ($\lambda=0.90$) leading to sharp increases in cost as $\alpha$ increases. This behavior confirms that high-risk aversion combined with high confidence levels results in the most conservative and expensive cost allocation strategies, whereas lower $\lambda$ values maintain cost stability across different $\alpha$ levels.
(a) OBJ vs. $\lambda$ for different $\alpha$
(b) $\mathbf{CVaR}$ vs. $\lambda$ for different $\alpha$
Figure 3: Relationship between $\lambda$, $\mathbf{OBJ}$, and $\mathbf{CVaR}$ for varying $\alpha$ levels.
Key Insights
From the analysis, the following critical insights emerge:
  • Balancing Cost and Risk: Increasing $\lambda$ improves resilience against extreme cost scenarios but does so at the expense of higher total expenditures. This trade-off is most pronounced for higher $\alpha$ values, where tail-risk mitigation is prioritized.
  • Diminishing Returns in CVaR Reduction: While higher $\lambda$ values significantly reduce CVaR initially, the rate of reduction slows beyond $\lambda=0.5$. This suggests that further increases in risk aversion do not necessarily translate into proportionate improvements in extreme risk mitigation.
  • Growing Cost Disparity Across $\alpha$ Levels: The widening gap in objective values for different $\alpha$ settings at higher $\lambda$ levels indicates that high-risk-averse settings amplify the sensitivity of the model to changes in confidence levels.
  • Non-Monotonic CVaR Behavior: At specific configurations, particularly $\lambda=0.8$ and $\alpha=0.90$, CVaR exhibits temporary stabilization before further decreasing at $\lambda=1.0$. This behavior suggests that the model optimally redistributes costs to balance both mean costs and tail risks dynamically.
Figure 4: Impact of $\alpha$ on the Objective Value for fixed $\lambda$ values.
Implications for Decision-Makers
These findings offer valuable insights for strategic decision-making:
  • For Risk-Averse Stakeholders (e.g., $\alpha=0.90$): Higher $\lambda$ values effectively minimize CVaR, making them suitable for industries prioritizing financial stability in volatile environments. However, this approach incurs significant additional costs.
  • For Cost-Sensitive Operators (e.g., $\alpha=0.25$): Lower $\lambda$ values ensure cost-efficient operations while maintaining moderate exposure to extreme risks, making them ideal for stakeholders with tighter budget constraints.
  • Optimal Risk Balancing Strategy: The presence of diminishing returns in CVaR reduction suggests that a moderate risk-aversion setting ($\lambda\approx 0.5$) may offer the best trade-off between cost efficiency and worst-case scenario protection.
In summary, the results emphasize the importance of balancing cost efficiency with resilience against extreme cost events. The widening gap in objective values across $\alpha$ levels suggests that models incorporating dynamic risk-preference adjustments can better align with diverse operational realities. Future work could explore how additional constraints, such as capacity limitations, supply fluctuations, and intermodal transfer restrictions, interact with risk-based cost optimizations.
5.3.2 Analysis of Cost Components
To gain a holistic understanding of the cost structure within the optimization framework, we analyze the aggregated cost components across all scenarios. The primary cost components include: Supply Cost, which represents the expenses associated with sourcing decisions; Transportation Cost, costs incurred in moving freight; Unmet Demand Penalty, which quantifies costs associated with unmet demand; and Emissions Penalty, reflecting the additional cost of exceeding predefined emissions limits. This analysis addresses Policy Question 1 by demonstrating how the model balances cost, risk, and decision-making under uncertainty, optimizing resource allocation to minimize costs while maintaining service levels and sustainability compliance. It also provides insights into Policy Question 2, showing how sustainability constraints affect cost distributions, and Policy Question 6, which examines the impact of capacity limitations on total costs and unmet demand.
Cost Drivers and Trade-offs in Optimization
The cost distribution, illustrated in Figure 5, is shaped by the model’s objective function and key operational constraints. The interplay between transportation, supply, unmet demand, and emissions penalties highlights the trade-offs involved in optimizing intermodal logistics.
  • Transportation Costs (58.1%) dominate total expenditures, reflecting the high logistical costs of intermodal freight movements. While the model prioritizes cost minimization, transportation remains the primary driver due to network topology and train capacity limits. To mitigate expenses, the model optimally assigns containers to the nearest supply locations, minimizing per-container transportation costs.
  • Supply Costs (34.8%) are influenced by sourcing decisions. Since supply nodes have finite capacities, the model balances cost-efficient procurement with constraints related to hub accessibility and demand fulfillment. Selecting lower-cost suppliers may increase emissions penalties or create mismatches between supply locations and demand hubs.
  • Unmet Demand Penalties (4.7%) arise from capacity constraints. The model strategically reduces unmet demand to minimize associated penalties, ensuring that available transportation capacity is used efficiently. However, limiting freight movement in low-demand scenarios may lower transportation costs while increasing unmet demand penalties.
  • Emissions Penalties (2.4%) remain relatively small, as the model integrates emissions constraints as a soft penalty rather than an absolute limit. While the current configuration allows minor exceedances, enforcing stricter sustainability targets would shift costs toward transportation and supply expenses.
The results underscore the interdependencies between cost components—minimizing one often leads to trade-offs
Figure 5: Aggregated cost breakdown across all scenarios.
Implications for Decision-Makers
The cost breakdown provides actionable insights for stakeholders:
  • Optimizing transportation operations is crucial, as it represents the largest cost component. Efficient route selection, train scheduling, and capacity planning can significantly reduce expenses by minimizing unnecessary freight movements and optimizing modal shifts. Additionally, reducing inefficiencies in transshipment operations and leveraging real-time data for dynamic scheduling can further cut costs. Stakeholders must also consider congestion effects at hubs, as delays and rerouting can indirectly increase transportation expenses.
  • Sustainability constraints have a minimal cost impact, reinforcing the model’s ability to balance emissions reduction with economic feasibility. The low emissions penalties suggest that the model effectively guides stakeholders toward environmentally responsible logistics without imposing excessive costs. Optimizing routes and sourcing locations inherently reduces emissions while maintaining cost efficiency, making the model a valuable tool for companies aiming to meet carbon reduction targets without sacrificing profitability.
  • Capacity constraints affect total costs, especially under demand uncertainty. While increasing capacity can reduce unmet demand penalties, the model demonstrates that dynamic allocation and smarter load balancing can achieve similar benefits without major infrastructure investments. Rather than expanding physical capacity, decision-makers should focus on adaptive planning, demand forecasting, and flexible routing to minimize service disruptions and cost escalations.
Broader Research Implications
This analysis highlights the importance of an integrated optimization approach, where transportation efficiency, sustainability, and capacity planning are addressed collectively rather than as independent factors. The results demonstrate the model’s effectiveness in supporting logistics providers, policymakers, and industry stakeholders in developing balanced, cost-efficient, and sustainable intermodal transportation strategies. Future research could explore the incorporation of dynamic pricing mechanisms to mitigate unmet demand penalties and evaluate alternative routing strategies to further optimize transportation costs. These findings provide a foundation for strategic policy development in intermodal freight planning, ensuring both economic viability and environmental sustainability in logistics operations.
5.3.3 Impact of Train Capacity Constraints on Demand and Total Cost
Train capacity constraints play a crucial role in intermodal transport, especially under stochastic demand and varying train capacity. To maintain computational efficiency, this experiment adopts a streamlined setup. The study examines a network of 4 freight trains operating under 8 demand scenarios across 22 intermodal hubs, with a total demand of 1013 containers, averaging 32 per train. This setup maintains the key stochastic elements while focusing on the analysis of the trade-offs between capacity expansion, cost reduction, and service reliability. The results, summarized in Table 3, shows how the capacity of the train influences unmet demand and costs, directly addressing the sixth policy question by identifying the threshold beyond which additional capacity offers diminishing cost savings.
Table 3: Impact of Train Capacity Constraints on Cost and Demand Satisfaction
Train Capacity Total Cost Unmet D D Met (%) Max Unmet D Min Unmet D Std Dev of Unmet D
4 249,400 453 55.28 18 10 3.10
5 214,015 293 71.08 14 5 2.61
6 182,709 178 82.43 13 1 3.64
7 156,209 70 93.09 13 0 2.93
8 137,284 13 98.72 3 0 0.90
9 127,593 9 99.11 2 0 0.67
10 121,681 9 99.11 2 0 0.67
Effect of Train Capacity on System Performance
Train capacity constraints directly impact cost and demand satisfaction, with increasing capacity leading to significant cost reductions and improved service levels. However, beyond a certain threshold, additional capacity yields diminishing returns. As shown in Table 3, increasing capacity from 4 to 8 containers per station significantly lowers costs and unmet demand, but further expansion to 10 containers results in only marginal improvements.
This trend is driven by the interaction between the train capacity constraint and the model’s objective:
  • Capacity Bottlenecks at Lower Levels: When capacity is low, the constraint (7) restricts container allocations, leading to unmet demand penalties. Increasing capacity alleviates these bottlenecks, improving system efficiency.
  • Scenario-Specific Demand Surges: With 8 demand scenarios, fluctuating demand levels create imbalances. At lower capacities, trains struggle to absorb peak demands, causing demand shortages. As capacity expands, the model gains more flexibility, reducing cost volatility.
  • Trade-offs in Cost Optimization: The model balances minimizing unmet demand penalties and optimizing transportation efficiency. While increasing capacity initially reduces penalties, beyond 8 containers per station, cost savings plateau as the system reaches near-optimal efficiency.
Interplay Between Cost and Unmet Demand
The relationship between total cost and unmet demand is further illustrated in Figure 6, which provides a visual representation of how capacity constraints shape system performance. At low capacity levels, the high unmet demand significantly contributes to the total system cost due to penalty terms associated with demand shortfalls. As capacity increases, unmet demand declines rapidly, leading to substantial cost reductions. However, beyond 8 containers per station, additional capacity improvements yield only marginal cost savings. This behavior suggests that while increasing capacity is effective in reducing costs initially, other factors such as supply constraints, competition for train slots, and network-wide operational limits begin to dominate decision-making at higher capacity levels.
The standard deviation of unmet demand across trains provides additional insights into how evenly shortages are distributed. At low capacity (4 containers per station), the system experiences significant demand imbalances, with the maximum unmet demand reaching 18 containers for a single train, while the minimum unmet demand remains 10 containers. This suggests that certain routes or stations experience significantly more congestion than others.
Figure 6: Total Cost and Unmet Demand vs. Train Capacity.
As capacity increases, unmet demand becomes more evenly distributed, with both its maximum value and standard deviation decreasing. By capacity 10, the maximum unmet demand is just 2 containers, and the standard deviation drops to 0.67, reflecting a more balanced use of available transport resources. This stabilization suggests that beyond a certain threshold, train capacity is no longer the primary bottleneck, other operational constraints, such as schedule synchronization, become more dominant in affecting system’s performance.
Key Insights
The following key takeaways emerge from the results:
  • Increasing train capacity significantly improves demand satisfaction, but beyond a certain threshold (8 containers per station in our case), further cost reductions are minimal due to other constraints becoming binding (e.g., supply availability and schedule alignment).
  • At lower capacities, demand imbalances are more pronounced, leading to high unmet demand variability across trains. This variability diminishes at higher capacities, indicating a more balanced network.
  • The model balances the trade-off between cost and demand fulfillment, minimizing penalties from unmet demand while ensuring train utilization remains efficient.
  • Network performance is highly sensitive to the available capacity per station, which is often influenced by external factors such as demand fluctuations and competition from other companies utilizing the same train services. Given these uncertainties, planning should focus on adapting to dynamic availability.
These insights reinforce the importance of strategic capacity planning in intermodal transport. As spot capacity availability increases, the system gains more flexibility in allocating containers, reducing the risk of unmet demand. However, relying solely on capacity expansion may not be the most cost-effective approach. Instead, decisions should prioritize adaptive scheduling, dynamic capacity allocation, and improved coordination among stakeholders to enhance system resilience and maximize efficiency under uncertain capacity conditions.
5.3.4 Sensitivity of Cost Metrics to Risk Preferences and Scenario Variability
A key challenge in stochastic intermodal transportation planning is ensuring decision robustness across cost-risk trade-offs and uncertainty scenarios. While prior analyses have examined the impact of $\lambda$ and $\alpha$ on total cost and CVaR, it is equally important to assess how these parameters influence cost variability and model stability. This section investigates, the divergence between objective value (OBJ) and expected total cost ($E(TC)$), cost variability across scenarios, and the stability of cost structures under different risk settings. Figures 7, 9, and 8 provide a visualization of these relationships.
Figure 7: Variation of OBJ Values with E(TC) for Different $\alpha$ Levels.
This analysis addresses the first policy question by illustrating how $\lambda$ and $\alpha$ affect cost predictability, resilience, and financial trade-offs. It also provides insights into the seventh policy question, highlighting at what point additional risk mitigation yields diminishing returns. Additionally, the findings contribute to the third policy question by demonstrating how different risk aversion levels impact cost robustness under varying demand conditions.
Discrepancy Between Objective Value and Expected Total Cost
A key observation from Figure 7 is the increasing gap between OBJ and $E(TC)$ as $\lambda$ rises, particularly at higher $\alpha$ values. At low $\lambda$, the two metrics remain closely aligned, indicating that when risk aversion is minimal, the model prioritizes cost efficiency with little regard for extreme scenarios. However, as $\lambda$ increases, OBJ becomes increasingly dominated by CVaR penalties, reflecting the fundamental trade-off between cost minimization and resilience to worst-case outcomes.
Scenario Variability and Cost Stability
Figure 8 shows that at low $\lambda$ values, total cost fluctuates significantly across scenarios, as the model optimizes for average cost rather than stability. As $\lambda$ increases, variability decreases, leading to more predictable cost structures. However, beyond $\lambda\approx 0.8$, further risk aversion yields diminishing reductions in cost fluctuations, suggesting that excessive risk mitigation comes at a steep cost with limited additional benefits.
Insights from CVaR Sensitivity Analysis
Figures 9 highlight how risk preferences influence extreme cost outcomes. While increasing $\lambda$ substantially reduces CVaR, the effect is more pronounced at higher confidence levels ($\alpha=0.90$) compared to lower ones. Notably, the gains from increasing $\lambda$ taper off beyond $\lambda\approx 0.5$, reinforcing that moderate risk aversion ($\lambda$ between 0.4-0.6) strikes an effective balance between mitigating extreme costs and maintaining overall cost efficiency.
Figure 8: Expected total cost values as a function of $\lambda$ and $\alpha$ illustrating the sensitivity of total cost to changes in $\lambda$ and $\alpha$.
Decision Robustness and Practical Implications
From a practical decision-making standpoint, these results reveal crucial insights into how risk preferences influence operational consistency and cost predictability:
  • For cost-minimization strategies ($\lambda\leq 0.3$): High cost variability across scenarios, more aggressive cost allocation, which may result in unanticipated financial risks when extreme-cost scenarios materialize, and suitable for stakeholders prioritizing efficiency over stability, particularly in low-volatility environments.
  • For balanced risk strategies ($\lambda\approx 0.5$): Moderate cost stability with significant CVaR reduction, lower variability in total costs, providing a predictable financial structure without excessive cost premiums, and suitable for stakeholders seeking a risk-aware but cost-efficient approach.
  • For highly risk-averse strategies ($\lambda\geq 0.8$): minimal cost fluctuations, leading to greater financial predictability, substantial increase in total cost, reducing cost efficiency in favor of extreme-risk protection, and best suited for industries with high exposure to financial uncertainty, such as global supply chains with high operational risk.
Figure 9: Visualization of CVaR values under varying $\lambda$ and $\alpha$ highlighting the sensitivity of CVaR to changes in $\lambda$ and $\alpha$.
Broader Implications and Future Considerations
These insights directly address the policy questions on cost-risk balancing and decision adaptability, demonstrating that the optimal choice of $\lambda$ and $\alpha$ depends on the decision-making priorities of different stakeholders. For future research, exploring interactions between train capacity constraints, emissions penalties, and risk-aversion strategies could further enhance the practical applicability of these findings. Additionally, expanding the scenario space to include real-world disruptions, such as terminal congestion and supply chain bottlenecks, would provide a more comprehensive understanding of risk-sensitive transportation planning.
5.3.5 Analyzing the Influence of Emissions Cap on Model Performance
The incorporation of emissions constraints in optimization models is a crucial factor in balancing cost efficiency with environmental sustainability. This subsection evaluates the impact of emissions caps ($\epsilon$) on the total cost structure, considering different risk-aversion levels ($\lambda$) and confidence levels ($\alpha$). The results, presented in Table 4 and illustrated in Figures LABEL:fig:OBJ_vs_E_.25, LABEL:fig:OBJ_vs_E_.5, LABEL:fig:OBJ_vs_E_.75, and LABEL:fig:OBJ_vs_E_.9, provide insights into how the emissions constraint interacts with other optimization components, shaping overall system performance.
This analysis directly addresses the second policy question on trade-offs between cost efficiency and sustainability, demonstrating how varying emissions caps impact the cost structure. Additionally, it provides insights into the first policy question by showing how decision-making under uncertainty is influenced by risk parameters when environmental constraints are introduced.
Impact of Emissions Cap on Objective Value.
Table 4 shows a clear trend: as the emissions cap increases, the objective value decreases across all configurations of $\lambda$ and $\alpha$. This behavior is a direct consequence of the constraint (9) which restricts emissions by penalizing excess emissions beyond $\epsilon$. A lower cap enforces stricter compliance, increasing the overall penalty and leading to higher costs. Conversely, a relaxed cap provides flexibility in route selection and transport decisions, reducing total costs. However, the objective function eventually stabilizes, indicating that beyond a certain point, emissions constraints cease to be the dominant cost driver.
Sensitivity to Risk-Aversion and Confidence Levels.
The model’s behavior under different risk settings highlights how emissions constraints interact with extreme cost mitigation. At higher values of $\lambda$, the optimization prioritizes risk reduction by minimizing tail-end cost scenarios, resulting in larger cost reductions when $\epsilon$ is relaxed. Similarly, higher $\alpha$ values amplify these effects, as the model becomes more sensitive to extreme cases in cost distribution. Figures LABEL:fig:OBJ_vs_E_.25LABEL:fig:OBJ_vs_E_.9 show that as risk aversion increases, the cost gap between different emissions caps widens. This reflects a trade-off: tight emissions constraints increase system resilience at the expense of higher costs.
Table 4: Objective Values for Different Emission Caps with Varying Lambda($\lambda$) and Alpha ($\alpha$) Values (scaled to $10^{3}$).
$\lambda=0.25$ $\lambda=0.50$ $\lambda=0.75$ $\lambda=0.90$
$\boldsymbol{\epsilon}$ $\alpha=0.25$ $\alpha=0.50$ $\alpha=0.70$ $\alpha=0.95$ $\alpha=0.25$ $\alpha=0.50$ $\alpha=0.70$ $\alpha=0.95$ $\alpha=0.25$ $\alpha=0.50$ $\alpha=0.70$ $\alpha=0.95$ $\alpha=0.25$ $\alpha=0.50$ $\alpha=0.70$ $\alpha=0.95$
25 251.7 261.9 267.2 278.1 258.4 278.3 288.3 309.2 264.5 293.8 307.3 338.6 268.3 302.3 317.6 354.4
50 247.6 257.9 263.2 274.3 254.4 274.3 284.4 304.9 260.6 289.6 303.9 334.2 264.4 298.4 313.9 350.7
75 244.0 254.6 259.6 270.8 251.0 271.1 281.0 301.8 257.4 287.4 301.3 331.9 261.5 296.4 312.7 349.4
100 240.0 250.9 255.6 266.5 246.9 267.7 277.3 298.2 253.9 283.8 297.8 328.8 257.9 293.6 310.2 346.8
125 238.4 248.7 254.3 264.1 245.4 266.1 276.4 296.4 252.0 282.4 297.7 327.5 255.7 292.0 309.6 345.3
150 235.1 245.2 250.5 260.7 241.8 262.9 272.8 293.1 249.0 280.4 294.9 325.4 253.3 290.8 308.1 344.2
175 233.4 244.5 250.1 259.8 240.5 262.0 272.3 292.1 247.6 279.0 294.4 323.8 251.8 289.8 307.6 343.2
200 230.9 241.6 247.1 256.9 238.1 259.5 270.2 289.8 245.3 277.3 293.0 322.7 249.6 287.9 306.5 341.9
225 230.4 241.4 246.9 256.5 237.7 259.1 270.0 288.9 245.2 277.3 293.0 321.1 249.5 288.0 306.5 340.4
250 228.7 239.6 245.1 254.6 235.9 257.6 268.5 287.4 243.1 275.5 291.5 320.4 247.4 286.2 305.2 339.8
275 228.3 239.1 244.6 254.1 235.4 257.1 268.0 287.1 242.6 275.5 291.0 319.4 246.9 285.7 304.7 338.8
300 227.5 238.4 243.9 253.3 234.7 256.4 267.3 286.0 241.9 274.4 290.4 318.4 246.2 285.1 304.1 337.8
325 228.0 238.4 243.9 253.3 235.2 256.9 267.3 286.0 242.4 274.4 290.4 318.9 246.2 285.6 304.6 337.8
350 227.5 238.4 243.9 253.3 234.7 256.4 267.3 286.0 241.9 274.4 290.4 318.4 246.2 285.1 304.1 337.8
375 227.5 238.4 243.9 253.3 234.7 256.4 267.3 286.0 241.9 274.4 290.4 318.4 246.2 285.1 304.1 337.8
Plateauing Effects and Non-Monotonic Trends.
The numerical results also exhibit plateauing behavior in objective values, particularly for higher $\epsilon$. This can be attributed to the fact that once emissions penalties no longer contribute significantly to cost, the model shifts focus to optimizing other components, such as transportation and supply costs. Furthermore, minor fluctuations in objective values at specific $\lambda$ and $\alpha$ combinations suggest potential threshold effects within the solver. These fluctuations indicate instances where the model finds alternative feasible solutions that marginally alter cost distribution but do not substantially impact overall performance.
Practical Implications for Sustainability and Cost Optimization.
The results emphasize that sustainability goals can be achieved without drastically increasing costs, provided emissions constraints are integrated strategically. At moderate $\lambda$ values, the model effectively balances emissions penalties with cost minimization, ensuring that environmental constraints do not become disproportionately restrictive. This suggests that decision-makers can set emissions limits that align with regulatory requirements without incurring excessive operational expenses.
Broader Research Implications.
This analysis provides valuable insights into the interplay between cost efficiency, emissions control, and risk preferences. It validates the model’s capacity to optimize intermodal transportation under policy-driven constraints and highlights the effectiveness of risk-adjusted decision-making in guiding sustainable logistics planning. Future research could explore dynamic emissions pricing mechanisms or the incorporation of carbon credit constraints to further enhance cost-emissions trade-offs.
Figure 10: Objective value versus emission cap ($\epsilon$) for different $\alpha$ values at $\lambda$ = 0.25, 0.50, 0.75, and 0.90.
5.3.6 Incorporating Intermodal Transfer Time and Cost
To enhance the applicability and realism of the proposed model, we extend the formulation to include intermodal transfer time and transfer cost, capturing the handling delays and financial expenditures incurred when transitioning goods from road to rail at intermodal hubs. While logistics companies cannot directly control these factors because they are managed by terminal operators and rail service providers their inclusion provides a clearer understanding of their impact on cost structure and scheduling decisions. Companies must ensure containers arrive at hubs with sufficient lead time to accommodate transfer activities and scheduled train departures.
Since the model is already designed to ensure timely container arrivals at intermodal hubs, additional experiments focusing solely on transfer components will not be carried out in this study. The key distinction in this extension lies in the addition of transfer costs per container and the time required for modal shifts. Although stakeholders cannot directly optimize intermodal transfers, incorporating these parameters strengthens the model’s applicability. Future research could build on this extension by modeling the entire transportation process from the supply location to final consumer, offering deeper insights into system-wide coordination across multiple hubs and transportation modes.
Motivation and New Parameters
Intermodal transfer activities such as unloading, sorting, and reloading introduce delays and costs that influence operational efficiency and scheduling. While not directly controllable by stakeholders in this study, these parameters provide critical insights into cost interplay and constraints within intermodal logistics. Their inclusion enhances the model’s predictive accuracy and practical applicability.
The following parameters are introduced to represent the effects of intermodal transfers:
  • $\tau_{ij}^{\text{tr}}$: Transfer time required to move goods from road to rail at intermodal station $j$, originating from location $i$. This parameter reflects delays from handling, processing, and equipment usage at the hub.
  • $C_{ij}^{\text{tr}}$: Transfer cost incurred for moving goods from road to rail at station $j$, originating from location $i$. This cost captures labor, equipment fees, and infrastructure charges associated with the intermodal transfer.
Enhanced Objective Function.
The inclusion of intermodal transfer cost $C_{\text{tr}}$ enhances the original objective function by explicitly capturing these additional costs. The revised objective function is as follows:
$\text{Min}\quad Z=$ $\sum_{i\in I}O_{i}y_{i}+(1-\lambda)\sum_{\omega\in\Omega}p_{\omega}\left(\sum_{i\in I}\sum_{j\in J}\sum_{n\in N}\sum_{t\in T}(C_{ij}+C^{\text{tr}}_{ij})x_{ijn\omega t}+\pi_{\omega}\sum_{n\in N}U_{n\omega}\right)$
$+\lambda\text{CVaR}+\sum_{t\in T}\rho\eta_{\omega}$ (26)
where $C_{ij}^{\text{tr}}$ accounts for the cost incurred during modal transitions at intermodal hubs. This modification ensures that transfer costs are evaluated alongside other operational components, allowing the model to account for all cost factors within intermodal logistics systems.
Updated Travel Time Constraints.
The introduction of transfer time $\tau_{ij}^{\text{tr}}$ requires adjustments to the travel time constraints. Goods transported from origin $i$ to station $j$ must now respect the total travel time, which includes both transit time $\tau(i,j)$ and transfer time $\tau_{ij}^{\text{tr}}$. The updated constraint is as follows:
$\sum_{i\in I}x_{ijn\omega t}\leq y_{i},\quad\forall j\in J,\,\forall n\in N,\,\forall\omega\in\Omega,\,\forall t\in T,\text{ if }t\geq\tau_{ij}+\tau_{ij}^{\text{tr}}.$ (27)
Additionally, goods cannot be transferred to the station earlier than the cumulative travel time allows, as enforced by the constraint below:
$\sum_{i\in I}x_{ijn\omega t}=0,\quad\forall j\in J,\,\forall n\in N,\,\forall\omega\in\Omega,\,\forall t\in T,\text{ if }t<\tau_{ij}+\tau_{ij}^{\text{tr}}.$ (28)
These constraints ensure that the total travel time, including intermodal transfers, aligns with the departure schedules of trains and prevents premature arrivals.
Stakeholder Implications and Future Extensions.
For logistics and manufacturing companies, intermodal transfer costs and delays are external factors determined by rail operators and hub administrators. While stakeholders cannot directly optimize these components, ensuring timely container arrivals at intermodal hubs allows the model to incorporate these constraints into the planning process effectively. Including intermodal transfer parameters lays the groundwork for future research on system-wide logistics optimization. A more comprehensive top-down approach could explicitly model transfer activities, capturing their interdependencies and impact across the broader supply chain.
In summary, while intermodal transfer costs and time remain fixed external inputs in the current model, their integration enhances the model’s realism and applicability. This extension not only strengthens its alignment with real-world constraints but also sets the stage for further investigations into intermodal coordination and optimization.
6 Conclusion and Future Work
This study introduces a novel two-stage stochastic optimization model for road-rail intermodal freight transportation, by integrating demand and capacity uncertainty, risk management, and sustainability objectives in a unified framework. Instead of focusing on isolated aspects of the intermodal network such as shipment allocation or carbon-efficient routing, our model uniquely combines these elements, offering a comprehensive approach to optimizing intermodal logistics under uncertainty. One of the key contributions of this study is the incorporation of real-time adaptive responses to stochastic elements such as train capacities and demand fluctuations while dynamically adjusting container allocation based on realized demand and capacity constraints. This adaptive approach ensures that decisions are aligned with operational realities, providing a level of flexibility and responsiveness that has not been previously achieved in intermodal freight optimization.
Another distinctive feature of our model is its dual emphasis on cost efficiency and environmental sustainability. Our model embeds carbon emission constraints directly into the optimization framework. By penalizing excess emissions and incorporating these penalties into the objective function, we ensure that carbon footprint reduction is a central consideration in the decision-making process. This approach aligns with regulatory frameworks, offering a practical solution for stakeholders seeking to comply with environmental standards. Furthermore, our model introduces a risk-averse optimization framework that leverages CVaR to manage extreme cost scenarios. While CVaR has been widely used in financial and supply chain contexts, its application to intermodal freight transportation under demand and capacity uncertainty is novel. Our findings demonstrate that the integration of CVaR not only enhances system resilience but also provides decision-makers with the flexibility to balance cost efficiency with risk aversion.
Building on these unique contributions, future research could explore several promising directions. First, the model could be extended to incorporate dynamic routing decisions, allowing for real-time adjustments to traffic conditions, infrastructure disruptions, and other unforeseen events. This would further enhance the model’s adaptability and applicability in dynamic logistics environments. Second, future work could investigate the integration of advanced predictive analytics and machine learning to improve demand forecasting and capacity planning. By leveraging real-time data and predictive models, the optimization framework could be further enhanced to anticipate and respond to fluctuations in demand and capacity more effectively. Finally, the model could be expanded to include multi-objective optimization techniques that simultaneously consider additional sustainability metrics, such as energy consumption, noise pollution, and social impacts. This would provide a more holistic approach to sustainable freight transportation, addressing a broader range of environmental and social concerns.
In conclusion, this study advances the field of intermodal freight transportation by addressing critical gaps in the literature and offering a robust, flexible, and sustainable optimization framework. The unique findings and contributions of this work provide a foundation for further exploration and innovation in the pursuit of efficient, resilient, and environmentally sustainable freight transportation systems.
Declaration of interests
The authors have no competing interests to declare that are relevant to the content of this article.
Appendix Appendix A Detailed Case Study Description
Appendix A.1 Network Description
The intermodal freight transport network for the case study used in this work consists of five major train routes connecting hubs across significant economic regions in the United States. These routes are designed to represent realistic operations in terms of distances and schedules and capacity constraints. Each hub serves as a critical transfer point for goods moving between road and rail modes, enabling efficient first-mile operations.
Appendix A.2 Example: Train A Route from Seattle to Chicago
Appendix A.2.1 Train Schedule
Train A operates on a meticulously planned schedule, departing from Seattle, WA, at 08:00 AM and arriving in Chicago, IL, at 06:00 PM on the third day. The route incorporates stops at key intermediate hubs, including Spokane, WA; Billings, MT; Fargo, ND; and Minneapolis, MN, each allowing for a six-hour dwell time. These dwell times are designed to accommodate essential freight handling operations, such as loading, unloading, and transfer activities, ensuring seamless flow throughout the network.
The detailed schedule for Train A is presented in Table A.1, capturing the precise arrival and departure times at each hub. This schedule serves as a foundation for analyzing transit times, operational efficiency, and the integration of first-mile and last-mile connectivity within the broader intermodal framework.
Appendix A.2.2 Hub-to-Hub Distances
The total distance for Train A is approximately 2,105 miles, with detailed distances between consecutive hubs provided in Table A.2. To ensure the highest level of accuracy, actual rail distances along the rail network were utilized. These distances reflect the practical realities of freight rail transport, accounting for deviations and curves inherent in rail lines compared to straight-line geodesic measurements.
By incorporating realistic rail distances, the analysis aligns closely with operational conditions, enabling precise calculations of transit times, fuel costs, and overall network efficiency. The detailed breakdown of distances between hubs in Table A.2 provides essential data for route optimization, further reinforcing the applicability of the study to real-world intermodal transport scenarios.
Appendix A.2.3 Nearby Facilities
Each intermodal hub along Train A’s route is strategically linked to nearby facilities, which act as critical supply or demand nodes for freight consolidation and distribution. These facilities, including warehouses and manufacturing plants, are selected based on their operational relevance and proximity to the hubs, ensuring seamless integration into the intermodal network.
Table A.3 provides a detailed overview of these facilities, including their names, road distances from the respective hubs, and geographical coordinates. The inclusion of these facilities not only highlights the importance of first-mile and last-mile connectivity in the intermodal system but also emphasizes the role of accurate facility-to-hub assignments in achieving cost-effective and timely freight operations. This linkage underscores the scalability and adaptability of the optimization framework in real-world applications.
Appendix A.3 Summary of All Train Routes
While Train A is discussed in detail to illustrate the methodology, the remaining four routes follow similar operational patterns, encompassing strategically selected hubs, defined schedules, and realistic distances. These routes represent a diverse set of intermodal freight connections across the United States, capturing variations in geographic coverage, transit times, and network complexity.
Table A.4 provides a comprehensive summary of the remaining routes, detailing their origins, destinations, intermediate hubs, total distances, and a high-level schedule overview. This summary serves as a concise yet informative reference, enabling practitioners and researchers to understand the broader applicability of the study. These additional routes enrich the case study by demonstrating the scalability of the proposed optimization framework to varying network configurations and operational scenarios.
Hub Departure Time Arrival Time Dwell Start Time Dwell End Time Travel Time (hours)
Buffalo, NY 12:30:00 18:57:00 19:27:00 24:57:00 6:27:00
Greencastle, PA 24:57:00 31:26:00 31:56:00 37:26:00 6:29:00
Columbus, OH 37:26:00 40:46:00 41:16:00 46:46:00 3:20:00
Cincinnati, OH 46:46:00 53:14:00 53:44:00 59:14:00 6:28:00
St. Louis, MO 59:14:00 62:24:00 62:54:00 68:24:00 3:10:00
Decatur, IL 68:24:00 75:30:00 76:00:00 81:30:00 7:06:00
Table A.1: Train A Schedule: Buffalo, NY, to Detroit, MI - Delray
Origin Hub Dest Hub Dist (m) Travel Time (hrs)
Buffalo, NY Greencastle, PA 280 6:27:00
Greencastle, PA Columbus, OH 540 6:29:00
Columbus, OH Cincinnati, OH 110 3:20:00
Cincinnati, OH St. Louis, MO 340 6:28:00
St. Louis, MO Decatur, IL 180 3:10:00
Decatur, IL Detroit, MI - Delray 310 7:06:00
Table A.2: Hub-to-Hub Distances and Transit Times for Train A
Appendix A.4 Model Parameters and Assumptions
This study employs realistic parameters for both road and rail transport to ensure accuracy and applicability. The following parameters and their ranges were used in the optimization framework:
Appendix A.4.1 Road Transport Parameters
The road segments connecting intermodal hubs to nearby facilities were modeled using an average speed of 55 miles per hour and a transportation cost of \$3.00 per mile. These values are derived from industry benchmarks for freight transport over road networks. The speed reflects variations in road types, traffic conditions, and loading/unloading delays, while the cost accounts for fuel expenses, maintenance, driver wages, and tolls. These parameters emphasize the role of facility proximity in minimizing first-mile and last-mile costs.
Appendix A.4.2 Penalty and Emission Parameters
To account for environmental and operational constraints in intermodal transport, the following parameters were included:
  • Emission Penalty ($\rho$): For exceeding emissions limits, penalties ranged from \$100 to \$500 per ton of CO2, reflecting the cost of environmental compliance and deterrence measures.
  • Unmet Demand Penalty ($\pi$): Costs for unmet demand ranged from \$1,000 to \$5,000 per container, capturing potential lost revenue and penalty fees for service failures.
  • Emission Caps ($\epsilon$): Over a 50-hour operational period, allowable emissions were capped at 50 to 150 tons of CO2, aligning with sustainability goals for the transport of 25–50 containers.
  • Carbon Emissions per Container ($E$): Emissions were estimated at 0.1–0.5 tons per container per hour, varying based on the transport mode and energy efficiency.
  • Container Preparation Costs ($O$): At origin facilities, preparation costs ranged from \$50 to \$200 per container, depending on facility efficiency and handling requirements.
Hub Facility Name Dist (m) Origin Coord (Lat, Lon) Dest Coord (Lat, Lon)
Buffalo, NY Fedex Warehouse 13 68 (43.04267661, -77.6750398) (42.8917322, -78.78303066)
Buffalo, NY Fedex Warehouse 14 16 (42.78052551, -78.84626101) (42.8917322, -78.78303066)
Greencastle, PA Fedex Warehouse 15 162 (41.2509818, -76.84433433) (39.76015595, -77.74182928)
Greencastle, PA Fedex Warehouse 16 190 (40.11206849, -74.97304753) (39.76015595, -77.74182928)
Columbus, OH Fedex Warehouse 17 23 (40.01861818, -82.9200998) (39.79411337, -82.95694498)
Columbus, OH Fedex Warehouse 18 91 (39.90979828, -84.18019515) (39.79411337, -82.95694498)
Cincinnati, OH Fedex Warehouse 19 10 (39.19316693, -84.47390985) (39.10476642, -84.53974175)
Cincinnati, OH Fedex Warehouse 20 65 (39.90979828, -84.18019515) (39.10476642, -84.53974175)
St. Louis, MO Fedex Warehouse 21 11 (38.69399342, -90.34175965) (38.68855031, -90.21324118)
St. Louis, MO Fedex Warehouse 22 29.4 (38.805481, -90.599304) (38.68855031, -90.21324118)
Decatur, IL Fedex Warehouse 23 186 (41.90215422, -87.64911463) (39.8736349, -88.91197339)
Decatur, IL Fedex Warehouse 24 79 (40.63198055, -89.45424028) (39.8736349, -88.91197339)
Detroit, MI - Delray Fedex Warehouse 25 19 (42.44998268, -83.1954107) (42.2961737, -83.12169962)
Detroit, MI - Delray Fedex Warehouse 26 28 (42.373589, -83.42067) (42.2961737, -83.12169962)
Table A.3: Nearby Facilities for Train A Route
Route Origin Destination Number of Intermediate Hubs Distance (miles)
Train B Taylor, PA New Orleans, LA 4 1,100
Train C Jacksonville, FL Ayer, MA 5 1,050
Train D Chicago 47th, IL Jacksonville, FL 4 1,200
Train E Dallas, TX Morrisville, PA 4 1,300
Table A.4: Summary of Additional Train Routes
Appendix A.4.3 Integration and Real-World Applicability
These parameters ensure a realistic representation of the trade-offs in intermodal freight transport. By incorporating penalties, emission constraints, and cost factors, the model balances operational efficiency with environmental sustainability. These considerations highlight the complexities of modern logistics, particularly in achieving carbon-neutral goals while maintaining service quality.
Appendix A.5 Reproducibility
This example provides sufficient detail to enable reproducibility and adaptation for similar studies. Researchers or practitioners can replicate the study by following these steps:
  1. 1.
    Train Schedules: Utilize detailed train schedules, including arrival, departure, and dwell times, as shown for Train A in Table A.1. These schedules ensure precise modeling of freight movements across hubs while accounting for operational constraints.
  2. 2.
    Hub-to-Hub Distances and Transit Times: Integrate hub-to-hub rail distances and corresponding transit times, as presented in Table A.2. These distances are either derived from actual rail line measurements or approximated using geodesic calculations to maintain spatial accuracy.
  3. 3.
    Facility Assignments: Assign nearby facilities to hubs based on proximity and road transport feasibility, as demonstrated in Table A.3. This ensures seamless integration within the network.
  • Train B: Taylor, PA → New Orleans, LA — NORFOLK INTERNATIONAL, VA, Charlotte, NC, Atlanta, GA, Huntsville, AL
  • Train C: Jacksonville, FL → Ayer, MA — Savannah, GA, Charleston, SC, Greensboro, NC, Pittsburgh, PA, Elizabeth Marine Terminal, NJ
  • Train D: Chicago 47th, IL → Jacksonville, FL — Louisville, KY, Georgetown, KY, Memphis, TN, Birmingham, AL
  • Train E: Dallas, TX → Morrisville, PA — New Orleans, LA, Birmingham, AL, Greer, SC, Norfolk Southern Bayview Yard, MD
While Train A serves as a detailed example in this study, summary details for the remaining four routes are also provided in the appendix (Table A.4). These summaries include key information on origins, destinations, intermediate hubs, and distances, allowing practitioners to extend the methodology to a broader intermodal network. By combining these datasets with the optimization framework described, this study ensures transparency and repeatability, laying the groundwork for future research and practical applications in intermodal freight transport optimization.